OpenStudy (anonymous):

I need to find the intercept: 5x + 2y = 10

OpenStudy (anonymous):

10+5x=2y

OpenStudy (anonymous):

which is it? Those are 2 different equations.

OpenStudy (anonymous):

The first one, I'm just trying to break it down

OpenStudy (anonymous):

Let's get it into slope intercept form (y = mx + b): 5x + 2y = 10 -5x -5x 2y = -5x + 10 /2 /2 /2 y = -(5/2)x + 5

OpenStudy (anonymous):

Why wouldn't it be 2y=10-5x?

OpenStudy (anonymous):

first off it'd be 10-5x=2y do you want both intercepts? To find the y intercept, it's the one that is b in the slope-intercept form y=mx+b So it'd be 5 (10/2) For the x intercept, set y to zero then solve for x -5=-(5/2)x 10=5x x is 2